OpenStudy (anonymous):

If f(x) = ax^2 + bx + a is divided by (x – b) a , b ≠ 0 , remainder is ‘a’ , then a is

OpenStudy (anonymous):

(A) b^2 (B) -1 (C) 0 (D) a + b

OpenStudy (turingtest):

I get -1... wanna see how?

OpenStudy (turingtest):

I'm not positive, but I don;t think I made any mistakes...

OpenStudy (anonymous):

Yes the answer is right, it matches the answer key.

OpenStudy (turingtest):

but I assume you would like to know how to do the problem, right?

OpenStudy (anonymous):

Yes

OpenStudy (turingtest):

here's I I did it (I'm going to have to use synthetic division, so bear with me because I have to draw it)...just wait

OpenStudy (anonymous):

Ok, no problem. Take your own time.

OpenStudy (turingtest):

now we are told that the remainder \[b^2(a+1)+a=a\to b^2(a+1)=0\]and since we are also told\[a,b\neq0\]that can only mean\[a+1=0\to a=-1\]